How do you change a problem such as, 2x^2=-6+8x, into a quadratic equation that can be solved using the quadratic formula?
Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

By moving all terms to one side:
2x^2=-6+8x
2x^2 - 8x =-6
2x^2 - 8x + 6 = 0
you could divide both sides by 2 to get:
x^2 - 4x + 3 = 0
Now, you can apply the "quadratic formula".
BUT in this case, it is easier to factor:
(x-1)(x-3) = 0
x = {1,3}

Answer by stanbon(75887)

2x^2=-6+8x,
------------------
Rearrange:
2x^2-8x+6 = 0
----
Divide thru by 2 to get:
x^2 - 4x + 3 = 0
------------------------
Use the quadratic formula:
x = [4 +- sqrt(16-4*3)]/2
----
x = [4 +- sqrt(4)]/2
----
x = [4 +- 2]/2
----
x = (6/2) or x = (2/2)
---
x = 3 or x = 1
=========================
Cheers,
Stan H.
==================